OpenStudy (anonymous):

Given that x=3sin@-2cos@ and y=3cos@+2sin@ , Show that x^2+y^2 is constant for all values of @. @=titre or whatever it spells . Hope you guys can help:)

hartnn (hartnn):

so did you square the 'x' ? what did u get ? x^2=(3sin@-2cos@ )^2 = ... ?

hartnn (hartnn):

you only need one relation to prove that : \(\huge \sin^2\theta+\cos^2 \theta =1\)

OpenStudy (anonymous):

okok thanks alot I've understood it!

hartnn (hartnn):

cool :) welcome ^_^